# Language Model Knowledge Distillation for Efficient Question Answering in Spanish

### 07 Dec 2023 v1

@@ 1 Introduction
In the last years, recent advancements in the field of NLP have panned the way for progress on a variety of downstream tasks, primarily by fine-tuning pre-trained language models (PLMs) (Vaswani et al., 2017; Devlin et al., 2019; Liu et al., 2019; Yang et al., 2019) on large-scale task-specific datasets. These models are typically trained on large and high-quality annotated corpora, which are are usually scarce for languages other than English, posing a significant disadvantage for progressing multilingual NLP research. Therefore, resources for languages such as Spanish, the fourth most spoken language, remain underrepresented. Moreover, despite the fact that some language models are available for the Spanish language (Gutiérrez-Fandiño et al., 2022; Cañete et al., 2020; de la Rosa et al., 2022; Pérez et al., 2022), operating them in resource-constrained environments remains an under-explored challenge. To overcome this limitation, we propose employing knowledge distillation (Jiao et al., 2020; Boreshban et al., 2021) to transfer the intrinsic knowledge of a large Spanish RoBERTa language model into a much compact and efficient model, named as SpanishTinyRoBERTa, while maintaining its performance. In particular, we focus on the question answering task (Rajpurkar et al., 2016; Carrino et al., 2019), where the goal of the model is to provide an answer given a context and a question over that context. Our experiments show that SpanishTinyRoBERTa attains negligible performance reduction on the question answering task with respect to the teacher model, whilst increasing significantly inference speed and reducing resource requirements. The ultimate goal of our work is to facilitate efforts in the development of efficient language models for the Spanish language, thereby breaking down computational barriers and promoting adoption of NLP technologies. Our training implementation and the source code is available at https://github.com/AdrianBZG/tinyroberta-distillation-qa-es.
@@ 3 Experimental Setup and Results
Model F1 (%) EM (%)
Spanish RoBERTa-large (teacher) 87.50 78.30
Multilingual BERT (mBERT) 77.60 61.80
Spanish BERT-base 82.15 73.59
Spanish RoBERTa-base 81.80 72.30
SpanishTinyRoBERTa (ours) 80.52 71.23
@@ A.2 Details on training hyperparameters and environment
Training was done using a single NVIDIA RTX A6000 GPU, with a batch size of 32, a learning rate of 3e-5, and maximum sequence length of 384 running for 20 epochs. For optimization purposes, mixed-precision training is employed, and gradient clipping with a max gradient norm of 1.0 is applied to improve training stability. We utilize the HuggingFace library (Wolf et al., 2020) for our training implementation and the source code is available at https://github.com/AdrianBZG/tinyroberta-distillation-qa-es.

### 16 Mar 2024 v2

@@ 1 Introduction
In the last years, recent advancements in the field of NLP have panned the way for progress on a variety of downstream tasks, primarily by fine-tuning pre-trained language models (PLMs) (Vaswani et al., 2017; Devlin et al., 2019; Liu et al., 2019; Yang et al., 2019) on large-scale task-specific datasets. These models are typically trained on large and high-quality annotated corpora, which are are usually scarce for languages other than English, posing a significant disadvantage for progressing multilingual NLP research. Therefore, resources for languages such as Spanish, the fourth most spoken language, remain underrepresented in terms of amount of training data available. Moreover, despite the fact that some language models are available for the Spanish language (Gutiérrez-Fandiño et al., 2022; Cañete et al., 2020; de la Rosa et al., 2022; Pérez et al., 2022), operating them in resource-constrained environments remains an under-explored challenge. To overcome this limitation, we propose employing knowledge distillation (Jiao et al., 2020; Boreshban et al., 2021) to transfer the intrinsic knowledge of a large Spanish RoBERTa language model into a much compact and efficient model, named as SpanishTinyRoBERTa, while maintaining its performance. In particular, we focus on the question answering task (Rajpurkar et al., 2016; Carrino et al., 2019), where the goal of the model is to provide an answer given a context and a question over that context. Our experiments show that SpanishTinyRoBERTa attains negligible performance reduction on the question answering task with respect to the teacher model, whilst increasing significantly inference speed and reducing resource requirements. The ultimate goal of our work is to facilitate efforts in the development of efficient language models for the Spanish language, thereby breaking down computational barriers and promoting adoption of NLP technologies.
@@ 3 Experimental Setup and Results
Model F1 (%) EM (%) Inference Speedup
Spanish RoBERTa-large (teacher) 87.50 78.30 1.0x
Multilingual BERT (mBERT) 77.60 61.80 3.0x
Spanish BERT-base 82.15 73.59 2.4x
Spanish RoBERTa-base 81.80 72.30 2.5x
SpanishTinyRoBERTa (ours) 80.52 71.23 4.2x
@@ 4 Conclusion
URM Statement
The authors acknowledge that at least one key author of this work meets the URM criteria of ICLR 2024 Tiny Papers Track.
@@ A.2 Details on training hyperparameters and environment
Training was done using a single NVIDIA RTX A6000 GPU, with a batch size of 32, a learning rate of 3e-5, and maximum sequence length of 384 running for 20 epochs. For optimization purposes, mixed-precision training is employed, and gradient clipping with a max gradient norm of 1.0 is applied to improve training stability. The distillation process took 5.3 hours to complete. We utilize the HuggingFace library (Wolf et al., 2020) for our training implementation and the source code is available at https://github.com/anonymous/anonymous.